OpenStudy (anonymous):

write the point-slop form of the equation of the line passing through the point (-2, 4) with a slope of -3 (1 point)

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -3 (-2,4)....x1 = -2 and y1 = 4 now we sub y - 4 = -3(x - (-2) y - 4 = -3(x + 2) <===point slope form any questions ?

OpenStudy (anonymous):

lol nope, could u help me with som more?

OpenStudy (texaschic101):

sure can...what ya got ?

OpenStudy (anonymous):

i need to find the value of x and y

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

sorry zeus...not so good at geometry

OpenStudy (anonymous):

@texaschic101 u have been helping me with it